OpenStudy (anonymous):

Explain how to find the sum 3+-8 using chips, a number line, or the rules for adding integers. Give the answer with your explanation.

OpenStudy (anonymous):

3 = (1+1+1) (3 chips) -8 = (1+1+1+1+1+1+1+1) (8 chips) 3 - 8 = (1+1+1) - (1+1+1+1+1+1+1+1) (3 chips - 8 chips) = (1+1+1) - 1 - 1 - 1 - 1 - 1 - 1 - 1 - 1 (distribute the negative sign) = 1 + 1 + 1 - 1 - 1 - 1 - 1 - 1 - 1 - 1 - 1 (make sure you do this correctly lol) = -5 (ta-da)

OpenStudy (anonymous):